1. Find F2–4 for the surfaces shown in Fig0.4 m 0.4m 0.6 m 3 2 4 1.2 m 2
Answer #1

Here length of surface 2 is not 0.4 m given. Let the length is x. So, 0.4+X+0.4-1.2 i.e. x-0.4 0. 0.6 m 3 2 4 1.2 m 2

Each surfaces (marked as 1,2,3,4) are rectangular.

So, Area of first surface (marked as 1)

Here, length is 0.4m and height is 0.6m.

So, area is 0.4*0.6=0.24 square meter.

Area of second surface is 0.4*0.6=0.24 square meter

Area of third surface is 0.4*0.6=0.24 square meter

Area of 4th surface (marked as 4)

(1/2)*1.2=0.6 square meter

Note: here question is confusing. F2 is not declared in the question. if F2 being the area of 2nd surface, which is marked as 2, then F2=0.24 square meter

Similarly, F3=0.24 square meter

and, F4=0.6 square meter